OpenStudy (anonymous):

determine the domain and range of the function f(x)=squareroot of x-11

OpenStudy (anonymous):

since you cannot take the square root of a negative number, the domain is found by solving \[x-11\geq 0\] for \(x\) (in one step)

OpenStudy (anonymous):

as for the range, \[\sqrt{\text{whatever}}\geq 0\] always